OpenStudy (anonymous):

solve equation by square root property r(2sq) + 36 = 0

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@XxNina26xX

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

it's written as r to the second power

jimthompson5910 (jim_thompson5910):

Hint: r^2 + 36 = 0 turns into r^2 = -36

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh wait, -6

OpenStudy (anonymous):

-18? sorry i'm graduating in a week from high school, i just need to finish this and i'm done but i need help :(

jimthompson5910 (jim_thompson5910):

have you learned about imaginary numbers? or complex numbers?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

well congratulation. But guessing isn't going to get you to graduation.

OpenStudy (anonymous):

i don't mean to guess, then help me please! lol

OpenStudy (anonymous):

do you understand square root property?

jimthompson5910 (jim_thompson5910):

if you start with r^2 and you replace r with -6, what do you get?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

-36!

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

- x - ?

OpenStudy (anonymous):

36

jimthompson5910 (jim_thompson5910):

r^2 = (-6)^2 (-6)^2 = (-6)*(-6) (-6)^2 = 36 ... notice how it's positive

jimthompson5910 (jim_thompson5910):

better

jimthompson5910 (jim_thompson5910):

you can take any number that's not 0 and square it it will always be positive

jimthompson5910 (jim_thompson5910):

so what this means is that r^2 = -36 has no real solutions but it does have solutions...they're just not considered real numbers

OpenStudy (anonymous):

oh so it would be -6?

jimthompson5910 (jim_thompson5910):

the convention is to make the letter i be equal to sqrt(-1) so i = sqrt(-1) notice how if you square both sides, you'll get i^2 = -1

jimthompson5910 (jim_thompson5910):

i = sqrt(-1) is known as an imaginary number

OpenStudy (anonymous):

okay :)

jimthompson5910 (jim_thompson5910):

technically all numbers are imaginary since we made them all up, so it's not really fair to call i = sqrt(-1) an imanigary number...but that's just how history played out

jimthompson5910 (jim_thompson5910):

imaginary*

OpenStudy (anonymous):

oh lol so i got 6 for r cause 6x 6 is 36

jimthompson5910 (jim_thompson5910):

anyways, i = sqrt(-1) so we'll use this to solve for r r^2 = -36 r = sqrt(-36) r = ??? ... how will we use the first equation above to simplify?

jimthompson5910 (jim_thompson5910):

6 is close...but not the full story

OpenStudy (anonymous):

6 x -1?

jimthompson5910 (jim_thompson5910):

I'll give you a further hint (if you're still stuck) r = sqrt(-36) r = sqrt(-1)*sqrt(36) r = ???

jimthompson5910 (jim_thompson5910):

and remember that i = sqrt(-1)

OpenStudy (anonymous):

-6

jimthompson5910 (jim_thompson5910):

sqrt(36) = ???

OpenStudy (anonymous):

so 36 x 36?

jimthompson5910 (jim_thompson5910):

i want you to start with r = sqrt(-1)*sqrt(36)

jimthompson5910 (jim_thompson5910):

and replace sqrt(-1) with ??? and also replace sqrt(36) with ???

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

what did we make sqrt(-1) equal to

OpenStudy (anonymous):

i?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

what is sqrt(36) equal to

OpenStudy (anonymous):

6?

jimthompson5910 (jim_thompson5910):

so r = sqrt(-1)*sqrt(36) turns into r = i*6 which flips to r = 6i

OpenStudy (anonymous):

why i?

jimthompson5910 (jim_thompson5910):

because i = sqrt(-1)

jimthompson5910 (jim_thompson5910):

notice how we just replaced sqrt(-1) with i

OpenStudy (anonymous):

but is i r?

jimthompson5910 (jim_thompson5910):

no i is its own thing

OpenStudy (anonymous):

oh.. okay thanks! here's another one

jimthompson5910 (jim_thompson5910):

wait, it turns out that -6i works as well (in much in the same way -6 works for r^2 = 36)

jimthompson5910 (jim_thompson5910):

so the two solutions to r^2 + 36 = 0 are r = 6i r = -6i